brobro19:

Graph each function and determine the y-intercept, then use the graph to determine the approximate value of the given expression. y = 3 Superscript x Baseline, 3 Superscript 1.2 a. 1; 3.7 c. 0; 3.9 b. 1; 2.9 d. 0; 3.5

lilianamendez2:

\(y = 3^2 \times _{ 3^2 1.2}\) ?

Vocaloid:

old question, answering to close y = 3^x is the function graph this function to estimate y = 3^1.2, find y = 1.2 on the graph and match it to the appropriate x-value